# Next-Generation Database Interfaces: A Survey of LLM-based Text-to-SQL

### v3

@@ Abstract
Generating accurate SQL from natural language questions (text-to-SQL) is a long-standing challenge due to the complexities in user question understanding, database schema comprehension, and SQL generation. Conventional text-to-SQL systems, comprising human engineering and deep neural networks, have made substantial progress. Subsequently, pre-trained language models (PLMs) have been developed and utilized for text-to-SQL tasks, achieving promising performance. As modern databases become more complex, the corresponding user questions also grow more challenging, causing PLMs with parameter constraints to produce incorrect SQL. This necessitates more sophisticated and tailored optimization methods, which, in turn, restricts the applications of PLM-based systems. Recently, large language models (LLMs) have demonstrated significant capabilities in natural language understanding as the model scale increases. Therefore, integrating LLM-based implementation can bring unique opportunities, improvements, and solutions to text-to-SQL research. In this survey, we present a comprehensive review of LLM-based text-to-SQL. Specifically, we propose a brief overview of the technical challenges and the evolutionary process of text-to-SQL. Then, we provide a detailed introduction to the datasets and metrics designed to evaluate text-to-SQL systems. After that, we present a systematic analysis of recent advances in LLM-based text-to-SQL. Finally, we discuss the remaining challenges in this field and propose expectations for future research directions.
@@ I Introduction
Figure 1: An example for LLM-based text-to-SQL selected from the BIRD dataset. A user proposes a question about football leagues. The LLM takes the question and the schema of its corresponding database as the input and then generates an SQL query as the output. The SQL query can be executed in the database and retrieve the content “The 5 leagues with the highest matches” to answer the user question.
@@
Text-to-SQL is a long-standing task in natural language processing research. It aims to convert (translate) natural language questions into database-executable SQL queries. Fig. 1 provides an example of a large language model-based (LLM-based) text-to-SQL system. Given a user question such as “Could you tell me the names of the 5 leagues with the highest matches of all time and how many matches were played in the said league?”, the LLM takes the question and its corresponding database schema as input and then generates an SQL query as output. This SQL query can be executed in the database to retrieve the relevant content to answer the user’s question. The above system builds a natural language interface to the database (NLIDB) with LLMs. Since SQL remains one of the most widely used programming languages, with over half (51.52%) of professional developers using SQL in their work, it is notable that only around a third (35.29%) of those developers are systematically trained [^1] https://survey.stackoverflow.co/2023, the NLIDB enables non-skilled users to access structured databases like professional database engineers [1, 2] and also accelerates human-computer interaction [3]. Furthermore, amid the research hotspot of LLMs, text-to-SQL can provide a potential solution to the prevalent hallucination [4, 5] issue by incorporating realistic content from the database to fill the knowledge gaps of LLMs [6]. The significant value and potential for text-to-SQL have triggered a range of studies on its integration and optimization with LLMs [7, 8, 9, 10]; consequently, LLM-based text-to-SQL remains a highly discussed research field within the NLP and database communities.
@@
Previous studies have made notable progress in implementing text-to-SQL and have undergone a long evolutionary process. Early efforts were mostly based on well-designed rules and templates [11], specifically suitable for simple database scenarios. In recent years, with the heavy labor costs [12] brought by rule-based methods and the growing complexity of database environments [13, 14, 15], designing a rule or template for each scenario has become increasingly difficult and impractical. The development of deep neural networks has advanced the progress of text-to-SQL [16, 17], which can automatically learn a mapping from the user question to its corresponding SQL [18, 19]. Subsequently, pre-trained language models (PLMs) with strong semantic parsing capacity have become the new paradigm for text-to-SQL systems [20], boosting their performance to a new level [21, 22, 23]. Incremental research on PLM-based optimization, such as table content encoding [24, 19, 25] and pre-training [20, 26], has further advanced this field. Most recently, the LLM-based approaches implementing text-to-SQL through in-context learning (ICL) [8] and fine-tuning (FT) [10] paradigm, reaching state-of-the-art accuracy with the well-designed framework and stronger comprehension capability compared to PLMs.
@@
The overall implementation details of LLM-based text-to-SQL can be divided into 3 aspects: 1. Question understanding: The NL question is a semantic representation of the user’s intention, which the corresponding generated SQL query is expected to align with; 2. Schema comprehension: The schema provides the table and column structure of the database, and the text-to-SQL system is required to identify the target components that match the user question; 3. SQL generation: This involves incorporating the above parsing and then predicting the correct syntax to generate executable SQL queries that can retrieve the required answer. The LLMs have proven to perform a good vanilla implementation [7, 27], benefiting from the more powerful semantic parsing capacity enabled by the richer training corpus [28, 29]. Further studies on enhancing the LLMs for question understanding [8, 9], schema comprehension [30, 31], and SQL generation [32] are being increasingly released.
Figure 2: A sketch of the evolutionary process for text-to-SQL research from the perspective of implementation paradigm. Each stage is presented with two representative implementation techniques. The timestamps for the stages are not exactly accurate; we set each timestamp according to the release time of the representative works of each implementation paradigm, with a margin of error of about one year before and after. The format is inspired from [29].
@@
augmented, leaf-datasets, text width=5.5em [BIRD [33], SQUALL [44], Spider-DK [39], modelnode-datasets, text width=43.5em] ] [ Cross-
lingual, leaf-datasets, text width=5.5em [DuSQL [34], CSpider [42], modelnode-datasets, text width=43.5em] ] [ Context-
dependent, leaf-datasets, text width=5.5em [CoSQL [35], Spider-SS [38], Spider-CG [38], SparC [43], modelnode-datasets, text width=43.5em] ] [ Robustness, leaf-datasets, text width=5.5em [ADVETA [37], Spider-SYN [40], Spider-Realistic [41], modelnode-datasets, text width=43.5em] ] ] ] [ Evaluation
@@
IV-A), leaf-methods, text width=6em [ Trivial
Prompt, leaf-methods, text width=6.5em [ Zero-shot, modelnode-methods, text width=4.5em [[7], [33], [27], [45], [46], [47], [48], [49], [9], [50], modelnode-methods,text width=36.2em]] [ Few-shot, modelnode-methods, text width=4.5em [[33], [8], [51], [52], [53], [54], [49], [9], [32], [55] , modelnode-methods,text width=36.2em]] ] [ Decomposition, leaf-methods, text width=6.5em [Coder-Reviewer [56], DIN-SQL [8], QDecomp [51], C3 [30], MAC-SQL [57], DEA-SQL [58],
SGU-SQL [32], MetaSQL [59], PET-SQL [60], PURPLE [61], modelnode-methods,text width=42.5em] ] [ Prompt
@@
ODIS [54], ACT-SQL [49], FUSED [65], DELLM [31], modelnode-methods,text width=42.5em] ] [ Reasoning
@@
SQL-CRAFT [55], FUXI [66], modelnode-methods,text width=42.5em] ] [ Execution
@@
DELLM [31], SQL-CRAFT [55], FUXI [66], PET-SQL [60], PURPLE [61], modelnode-methods,text width=42.5em] ] ] [ Fine-tuning
@@
IV-B), leaf-methods,text width=6em [ Vanilla
Fine-tuning, leaf-methods, text width=6.5em [[45], [9], [50], [53], modelnode-methods,text width=42.5em] ] [ Enhanced
@@
Augmentation, leaf-methods, text width=6.5em [DAIL-SQL [9], Symbol-LLM [50], CodeS [10], StructLM [70], modelnode-methods,text width=42.5em] ] [ Decomposition, leaf-methods, text width=6.5em [DTS-SQL [71], modelnode-methods,text width=42.5em] ] ] ] ]
Figure 3: Taxonomy tree of the research in LLM-based text-to-SQL. The display order in each node is organized by the released time. The format is adapted from [72].
@@
Despite the significant progress made in text-to-SQL research, several challenges remain that hinder the development of robust and generalized text-to-SQL systems [73]. Related works in recent years have surveyed the text-to-SQL systems in deep learning approaches and provided insights into the previous deep neural network and PLM-based research [74, 2, 29]. In this survey, we aim to catch up with the recent advances and provide a comprehensive review of the current state-of-the-art models and approaches in LLM-based text-to-SQL. We begin by introducing the fundamental concepts and challenges associated with text-to-SQL, highlighting the importance of this task in various domains. We then delve into the evolution of the implementation paradigm for text-to-SQL systems, discussing the key advancements and breakthroughs in this field. After the overview, we provide a detailed introduction and analysis of the recent advances in text-to-SQL integrating LLMs. Specifically, the body of our survey covers a range of contents related to LLM-based text-to-SQL, including:
@@
  • Datasets and Benchmarks: We provide a detailed introduction to the commonly used datasets and benchmarks for evaluating LLM-based text-to-SQL systems. We discuss their characteristics, complexity, and the challenges they pose for text-to-SQL development and evaluation.
  • Evaluation Metrics: We present the evaluation metrics used to assess the performance of LLM-based text-to-SQL systems, including both content matching-based and execution-based paradigms. We then briefly introduce the characteristics of each metric.
@@
  • Methods and Models: We present a systematic analysis of the different methods and models employed for LLM-based text-to-SQL, including in-context learning and fine-tuning-based paradigms. We discuss their implementation details, strengths, and adaptations specific to the text-to-SQL task from various implementation perspectives.
@@
  • Expectations and Future Directions: We discuss the remaining challenges and limitations of LLM-based text-to-SQL, such as real-world robustness, computational efficiency, data privacy, and extensions. We also outline potential future research directions and opportunities for improvement and optimization.
We hope this survey provides a clear overview of recent studies and inspires future research. Fig. 3 shows a taxonomy tree that summarizes the structure and contents of our survey.
@@ II Overview
Text-to-SQL is a task that aims to convert natural language questions into corresponding SQL queries that can be executed in a relational database. Formally, given a user question $\mathcal{Q}$ (also known as a user query, natural language question, etc.) and a database schema $\mathcal{S}$, the goal of the task is to generate an SQL query $Y$ that retrieves the required content from the database to answer the user question. Text-to-SQL has the potential to democratize access to data by allowing users to interact with databases using natural language without the need for specialized knowledge of SQL programming [75]. This can benefit various domains, such as business intelligence, customer support, and scientific research, by enabling non-skilled users to easily retrieve target content from databases and facilitating more efficient data analysis.
@@ II-A Challenges in Text-to-SQL
The technical challenges for text-to-SQL implementations can be summarized as follows:
@@ II-A1 Linguistic Complexity and Ambiguity
Natural language questions often contain complex linguistic representations, such as nested clauses, coreferences, and ellipses, which make it challenging to map them accurately to the corresponding part of SQL queries [41]. Additionally, natural language is inherently ambiguous, with multiple possible representations for a given user question [76, 77]. Resolving these ambiguities and understanding the intent behind the user question requires deep natural language understanding and the capability to incorporate context and domain knowledge [33].
@@ II-A2 Schema Understanding and Representation
To generate accurate SQL queries, text-to-SQL systems need to have a comprehensive understanding of the database schema, including table names, column names, and relationships between various tables. However, the database schema can be complex and vary significantly across different domains [13]. Representing and encoding the schema information in a way that can be effectively utilized by the text-to-SQL model is a challenging task.
@@ II-A3 Rare and Complex SQL Operations
Some SQL queries involve rare or complex operations and syntax in challenging scenarios, such as nested sub-queries, outer joins, and window functions. These operations are less frequent in the training data and pose challenges for text-to-SQL systems to generate accurately. Designing models that can be generalized to a wide range of SQL operations, including rare and complex scenarios, is an essential consideration.
@@ II-A4 Cross-Domain Generalization
Text-to-SQL systems often struggle to generalize across various database scenarios and domains. Models trained on a specific domain may not perform well on the proposed questions from other domains due to the variety in vocabulary, database schema structure, and question patterns. Developing systems that can effectively generalized to new domains with minimal domain-specific training data or fine-tuning adaptation is a significant challenge [78].
@@ II-B1 Rule-based Methods
Early text-to-SQL systems relied heavily on rule-based methods [11, 12, 26], where manually crafted rules and heuristics were used to map natural language questions to SQL queries. These approaches often involved extensive feature engineering and domain-specific knowledge. While rule-based methods achieved success in specific simple domains, they lacked the flexibility and generalization capabilities needed to handle diverse and complex questions.

### v4

@@ Abstract
Generating accurate SQL from users’ natural language questions (text-to-SQL) remains a long-standing challenge due to the complexities involved in user question understanding, database schema comprehension, and SQL generation. Traditional text-to-SQL systems, which combine human engineering and deep neural networks, have made significant progress. Subsequently, pre-trained language models (PLMs) have been developed for text-to-SQL tasks, achieving promising results. However, as modern databases and user questions grow more complex, PLMs with a limited parameter size often produce incorrect SQL. This necessitates more sophisticated and tailored optimization methods, which restricts the application of PLM-based systems. Recently, large language models (LLMs) have shown significant capabilities in natural language understanding as model scale increases. Thus, integrating LLM-based solutions can bring unique opportunities, improvements, and solutions to text-to-SQL research. In this survey, we provide a comprehensive review of existing LLM-based text-to-SQL studies. Specifically, we offer a brief overview of the technical challenges and evolutionary process of text-to-SQL. Next, we introduce the datasets and metrics designed to evaluate text-to-SQL systems. Subsequently, we present a systematic analysis of recent advances in LLM-based text-to-SQL. Finally, we make a summarization and discuss the remaining challenges in this field and suggest expectations for future research directions.
@@ I Introduction
Figure 1: An example of LLM-based text-to-SQL is selected from the BIRD [1] dataset. A user asks a question about football leagues. The LLM takes this question along with the schema of the corresponding database as input and generates an SQL query as output. The generated SQL query can be executed in the database, retrieves the content “The 5 leagues with the highest matches”, providing the answer to the user’s question.
@@
Text-to-SQL is a long-standing task in natural language processing, aiming to convert natural language (NL) questions into database-executable SQL queries. Fig. 1 illustrates an example of a large language model-based (LLM-based) text-to-SQL system. Given a user question such as “Could you tell me the names of the 5 leagues with the highest matches of all time and how many matches were played in the said league?”, the LLM takes the question and its corresponding database schema as input to generate an SQL query. Executing this SQL query in the database retrieves the relevant content needed to answer the user’s question. This system serves as a natural language interface to the database (NLIDB) using LLMs. Since SQL is one of the most widely used programming languages, with over half (51.52%) of professional developers employing it, but only around a third (35.29%) are systematically trained, [^1] https://survey.stackoverflow.co/2023 the NLIDB allows non-skilled users to access structured databases like professional database engineers [2, 3] and enhances human-computer interaction [4]. Furthermore, amid the research hotspot of LLMs, text-to-SQL can potentially mitigate prevalent hallucination issues by integrating realistic content from databases to fill LLMs’ knowledge gaps [5]. The significant value of text-to-SQL have triggered a range of studies on its integration and optimization with LLMs [6, 7, 8]; consequently, LLM-based text-to-SQL remains a highly discussed research field in the NLP and database communities.
@@
Previous traditional methods have made notable progress in implementing text-to-SQL. As shown in Fig. 2, these implementations have undergone a long evolutionary process. Early efforts were mostly based on well-designed rules and templates [9], specifically suitable for simple database scenarios. In recent years, due to the heavy labor costs [10] associated with rule-based methods and the growing complexity of database environments [11, 12, 13], designing a rule or template for each scenario has become increasingly difficult and impractical. The development of deep neural networks has advanced the progress of text-to-SQL [14, 15], which can automatically learn a mapping from the user question to its corresponding SQL [16, 17]. Subsequently, pre-trained language models (PLMs) with strong semantic parsing capabilities have become the new paradigm for text-to-SQL systems [18], boosting their performance to a new level [19, 20, 21]. Incremental research on PLM-based optimization, such as table content encoding [18, 17, 22] and pre-training [18, 23], has further advanced this field. Most recently, LLM-based approaches have implemented text-to-SQL through the in-context learning [7] and fine-tuning [24] paradigms, reaching state-of-the-art accuracy with well-designed frameworks and stronger comprehension capabilities compared to PLM-based methods.
@@
The overall implementation details of LLM-based text-to-SQL can be divided into three aspects: 1. Question understanding: The NL question is a semantic representation of the user’s intention, which the corresponding generated SQL query is expected to align with; 2. Schema comprehension: The schema provides the table and column structure of the database, and the text-to-SQL system is required to identify the target components that match the user question; 3. SQL generation: This involves incorporating the above parsing and then predicting the correct syntax to generate executable SQL queries that can retrieve the required answer. LLMs have proven to perform a successful vanilla implementation [6, 25], benefiting from the powerful semantic parsing capacity enabled by rich training corpus [26, 27]. Further studies on enhancing LLMs for question understanding [7, 8], schema comprehension [28, 29], and SQL generation [30] are being increasingly released.
Figure 2: A sketch of text-to-SQL research evolution outlines the advancement of implementation paradigms, each represented with different color backgrounds. Before 2023, the focus is on a selection of representative traditional studies. However, from 2023 onward, the emphasis shifts to the rapid advancements driven by LLMs, marking a significant acceleration in the field.
@@
augmented, leaf-datasets, text width=5.5em [BIRD [1], SQUALL [41], Spider-DK [36], BIRD-CRITIC [1], Spider 2.0 [42], modelnode-datasets, text width=43.5em] ] [ Cross-
lingual, leaf-datasets, text width=5.5em [DuSQL [31], CSpider [39], Spider-Vietnamese [43], modelnode-datasets, text width=43.5em] ] [ Context-
dependent, leaf-datasets, text width=5.5em [CoSQL [32], Spider-SS [35], Spider-CG [35], SparC [40], modelnode-datasets, text width=43.5em] ] [ Robustness, leaf-datasets, text width=5.5em [ADVETA [34], Spider-SYN [37], Spider-Realistic [38], Dr. Spider [44], modelnode-datasets, text width=43.5em] ] [ Long-context, leaf-datasets, text width=5.5em [BIRD-CRITIC [1], Spider 2.0 [42], modelnode-datasets, text width=43.5em] ] [ Specialized-
domain, leaf-datasets, text width=5.5em [BULL [45], modelnode-datasets, text width=43.5em] ] ] ] [ Evaluation
@@
IV-A), leaf-methods, text width=6em [ Decomposition, leaf-methods, text width=6.5em [Coder-Reviewer [46], DIN-SQL [7], QDecomp [47], C3 [28], MAC-SQL [48], DEA-SQL [49],
SGU-SQL [30], MetaSQL [50], $\textit{{R}}^{3}$ [51], PET-SQL [52], PURPLE [53], TA-SQL [54], MCS-SQL [55],
CHESS [56], SuperSQL [57], Distillery [58], RSL-SQL [59], Gen-SQL [60], ReFoRCE [61], modelnode-methods,text width=42.5em] ] [ Prompt
@@
ODIS [66], ACT-SQL [67], FUSED [68], DELLM [69], Dubo-SQL [70], TA-SQL [54],
MCS-SQL [55], SuperSQL [57], Distillery [58], RSL-SQL [59], ICRL [71], Gen-SQL [60],
SAFE-SQL [72], ReFoRCE [61], modelnode-methods,text width=42.5em] ] [ Reasoning
@@
SQL-CRAFT [74], FUXI [75], SuperSQL [57], CHASE-SQL [76], ReFoRCE [61], modelnode-methods,text width=42.5em] ] [ Execution
@@
DELLM [69], SQL-CRAFT [74], FUXI [75], $\textit{{R}}^{3}$ [51], PET-SQL [52], PURPLE [53], Dubo-SQL [70],
MCS-SQL [55], CHESS [56], Distillery [58], E-SQL [80], CHASE-SQL [76], RSL-SQL [59], modelnode-methods,text width=42.5em] ] ] [ Fine-tuning
@@
IV-B), leaf-methods,text width=6em [ Enhanced
@@
Augmentation, leaf-methods, text width=6.5em [DAIL-SQL [8], Symbol-LLM [82], CodeS [24], StructLM [83], Dubo-SQL [70], Distillery [58],
XiYan-SQL [84], modelnode-methods,text width=42.5em] ] [ Multi-task
Tuning, leaf-methods, text width=6.5em [SQL-LLaMA [85], DTS-SQL [86], KaSLA [29], MSc-SQL [87], XiYan-SQL [84], ROUTE [88], modelnode-methods,text width=42.5em] ] ] ] ]
Figure 3: Taxonomy tree of research in LLM-based text-to-SQL, taking inspiration from [89]; nodes are ordered by release time.
@@
Despite the significant progress made in text-to-SQL research, several challenges remain that hinder the development of robust and generalized text-to-SQL systems [90]. Recent related works have surveyed text-to-SQL systems using deep learning approaches and provided insights into previous deep neural network and PLM-based research [91, 3, 27, 92]. In this survey, we aim to catch up with recent advances and provide a comprehensive review of the current state-of-the-art models and approaches in LLM-based text-to-SQL. We begin by introducing the fundamental concepts and challenges associated with text-to-SQL, highlighting the importance of this task in various domains. We then delve into the evolution of the implementation paradigm for text-to-SQL systems, discussing the key advancements and breakthroughs in this field. After the overview, we provide a detailed introduction and analysis of the recent advances in text-to-SQL integrating LLMs. Specifically, the body of our survey covers a range of topics related to LLM-based text-to-SQL, including:
@@
  • Benchmarks and Metrics: We introduce the commonly used datasets and benchmarks for evaluating LLM-based text-to-SQL systems. We discuss their characteristics, complexity, and the challenges they pose for development and evaluation, along with the evaluation metrics.
@@
  • Methods: We systematically analyze different methods proposed for LLM-based text-to-SQL, including in-context learning and fine-tuning paradigms. We explore their implementation details, strengths, and specific adaptations, organized through well-designed categorization.
@@
  • Summarization and Expectations: We summarize the advantages and limitations of both LLM-based and traditional methods, as well as within LLM-based methods themselves. We discuss remaining challenges such as real-world robustness, computational efficiency, data privacy, and possible extensions. Additionally, we outline future research directions and opportunities for improvement.
To summarize our technical contributions, at the time of this paper’s release, we present the first LLM-based text-to-SQL survey in the community, establishing a systematic taxonomy for LLM-based text-to-SQL methods. This categorization provides analysis from technical implementation perspectives, clearly indicating their contributions to generating better SQL. Then, we further integrate insights while analyzing the trade-offs between LLM-based and traditional methods, as well as between in-context learning and fine-tuning paradigms for future developments. Finally, we highlight the remaining challenges in robustness, real-world deployment, efficiency, and domain adaptation, offering actionable directions for future research. Fig. 3 shows a taxonomy tree that summarizes the structure and contents of our survey.
@@ II Overview
Text-to-SQL is a task that aims to convert NL questions into SQL queries that can be executed in a relational database. It has the potential to democratize data access by allowing users to interact with databases using natural language, thus eliminating the need for specialized SQL programming knowledge [93]. This capability can greatly benefit domains such as business intelligence, customer support, and scientific research, by enabling non-specialized users to easily retrieve target content from databases and facilitating more efficient data analysis.
@@ II-A Challenges in Text-to-SQL
The technical challenges are summarized as follows:
@@ II-A1 Linguistic Complexity and Ambiguity
Natural language questions often contain complex linguistic representations, such as nested clauses, coreferences, and ellipses, which make it challenging to map them accurately to the corresponding parts of SQL queries [38]. Additionally, natural language is inherently ambiguous, with multiple possible representations for a given user question [94, 95]. Resolving these ambiguities and understanding the intent behind the user question requires deep natural language understanding and the capability to incorporate context and domain knowledge [1].
@@ II-A2 Schema Understanding and Representation
To generate accurate SQL queries, text-to-SQL systems must have a comprehensive understanding of the database schema, including table names, column names, and relationships between various tables. Representing and encoding schema information is challenging because the database schema is often complex and varies significantly across different domains [11].
@@ II-A3 Rare and Complex SQL Operations
Some SQL queries involve rare or complex operations and syntax in challenging scenarios, such as nested sub-queries, outer joins, and window functions. These operations are less frequent in the training data and pose challenges for text-to-SQL systems to generate accurately. Designing models that can generalize to a wide range of SQL operations, including rare and complex scenarios, is an essential consideration.
@@ II-A4 Cross-Domain Generalization
Text-to-SQL systems often struggle to generalize across various database scenarios and domains. Models trained on a specific domain may not perform well on questions from other domains due to variations in vocabulary, database schema structure, and question patterns. Developing systems that can effectively generalize to new domains with minimal domain-specific training data or fine-tuning adaptation is a significant challenge [96].
@@ II-B1 Rule-based Methods
Early text-to-SQL systems relied heavily on rule-based methods [9, 10, 23], where manually crafted rules and heuristics were used to map natural language questions to SQL queries. Rule-based systems excel at generating syntactically correct SQL queries by relying on predefined templates and strict grammatical rules, ensuring adherence to SQL syntax and structure. Their deterministic nature guarantees consistent outputs for specific input patterns, minimizing variability in query generation. However, these methods struggle to interpret linguistically complex or ambiguous natural language questions, such as those involving nested clauses, coreferences, or ellipses, and often fail to map these elements to database schema components like tables or columns. Their reliance on manual rules limits generalization to unseen question patterns or schema structures, requiring frequent updates for new domains. Additionally, encoding complex schemas with multi-table relationships into usable rules is labor-intensive and prone to errors.